A ball is thrown from the top of a building with a horizontal velocity of 11.9 m/s and takes 3.54 s to hit the ground. How far above the ground, in m, is the ball after 1.47 s of falling?
1
Expert's answer
2019-10-30T12:30:57-0400

The total height:


"H=\frac{gT^2}{2}"

The height of the ball after 1.47 s of falling:


"h=H-\frac{gt^2}{2}=\frac{gT^2}{2}-\frac{gt^2}{2}"

"h=0.5g(T^2-t^2)=0.5(9.8)(3.54^2-1.47^2)=50.8\ m"